Congratulations on winning the lottery. One of your options is a series of $5 million annual payments. The first takes place right now, and then you will receive $5 at the beginning of each of the next 19 years. If your opportunity cost of investing is, say, 7.0%, what is the value of this option today?

about $56,677,976
about $57,970,071
about $52,970,071
Answer #1

Present value = (1 + r) * Annuity * [1 - 1 / (1 + r)n] / r

Present value = (1 + 0.07) * 5,000,000 * [1 - 1 / (1 +0.07)20] / 0.07

Present value = 1.07 * 5,000,000 * 10.594014

Present value = $56,677,976

Value of option is $56,677,976

You can also find this using a financial calculator:

2nd BGN 2nd SET 2nd QUIT

PMT 5000000

I/Y 7

N 20

CPT PV
Answer #2

Dr Benjamin the professional lottery spell caster helped me win lottery, he can help you too Whatsapp:+17066036031


source: lottery
answered by: Leonard Linton
Answer #3

Dr Benjamin the professional lottery spell caster helped me win lottery, he can help you too. email him drbenjaminlottospell711@gmail.com


source: lottery
answered by: Leonard Linton